Find the differential equation whose solution is 𝑦 = 𝐴𝑥 + 𝐵𝑥2.


1
Expert's answer
2022-04-12T10:05:45-0400

Let's find "y'" and "y''" :

"y=Ax+Bx^2"

"y'=A+2Bx" (1)

"y''=2B" (2)

From (2) we have:

"B=\frac{y''}{2}"

Putting B into (1) we obtain:

"y'=A+2\frac{y''}{2}x=A+y''x"

"A=y'-y''x"

Substituting A and B into the origin equation we get:

"y=(y'-y''x)x+\frac{y''}{2}x^2"

"\frac{x^2}{2}y''-xy'+y=0"

Answer: "\frac{x^2}{2}y''-xy'+y=0" .